The following information exists for ABC Company:
Selling price per unit: $30    Variable expenses per unit: $21
Fixed expenses for the period: $60,000
Sales volume in units: 10,000

1. If advertising of $15,000 is spent to increase sales volume by 2,000 units,
operating income will increase by ??


2. Based on the information given above, ABC Company's contribution margin ratio will be ??
------------------------------------------------------------------------------------------

The following information exists for ABC Company:
Selling price per unit = $60
Variable expenses per unit = $45

If ABC's break-even sales revenue is $150,000 and sales revenue for April totals $140,000,
3. then for April, the company's operating loss will be ??
------------------------------------------------------------------------------------------

The following information exists for ABC Company:
Selling price per unit = $60      
Variable expenses per unit = $40
If ABC's breakeven point is 5,000 units and it sells 5,750 units in March,
4. its operating income will be ??

------------------------------------------------------------------------------------------

The following information exists for ABC Company:
Sales revenue: $300,000
Contribution margin ratio: 30%
Fixed expenses for the period: $60,000
Sales volume in units: 10,000

5. If sales revenue increases by $20,000, operating income will increase by ??
------------------------------------------------------------------------------------------

The following information exists for a firm:
Selling price per unit = $50
Contribution margin ratio = 30%
Fixed expenses per month = $30,000
Desired operating income = $18,000

6. The total revenues necessary to earn the desired operating income is ??

------------------------------------------------------------------------------------------

{Please show calculations for all questions. I been trying for a while and can't for the life of me figure this out! TIA!}
Answer #1

1) Total Contribution = (30-21)*10000 = 90000

Operating income = 90000-60000 = 30000

Increase in sales by 2000 units then,

Total contribution = (30-21)*12000 = 108000

Operating income = 108000-60000-15000 = 33000

Hence operating income will increase by 33000-30000 = $3000

2) Contribution margin ratio = (30-21)/30 = 30%

3) Contribution margin ratio = (60-45)/60 = 25%

Fixed costs = Break even sales * Contribution margin ratio = 150000*25% = 37500

Operating loss = (140000*25%) - 37500 = -2500

4) Contribution margin ratio = (60-40)/60 = 33.33%

Break even sales = 60*5000 = 300000

Fixed cost = 300000*33.33% = 100000

Operating income = { 5750*(60-40) } - 100000 = 15000

5) Selling price per unit = 300000/10000 = 30

Operating income at present = (300000*30%)-60000 = 30000

If sales revenue increases by $20000 then,

Operating income = (320000*30%)-60000 = 36000

Increase in operating income : 36000-30000 = 6000

6) Revenue required to earn desired operating income = (30000+18000)/30% = $160000